OpenStudy (anonymous):

Complete the square, rewrite y = 3x2 - 24x + 10 in general form.

OpenStudy (anonymous):

3x^2 - 24 = y - 10 3 (x^2 - 8) = y - 10 x^2 - 8 = (y-10)/3 x^2 - 8 + 16 = (y-10)/3 + 16 (x-4)^2 = (y-10)/3 + 16 (x-4)^2 - 16 = (y-10)/3 3(x-4)^2 - 54 = y - 10 y = 3(x-4)^2 - 44

OpenStudy (anonymous):

How did you get 16?

OpenStudy (anonymous):

@moneybird